# Learning Distances from Data with Normalizing Flows and Score Matching

### v1

@@ 4.3 Normalizing flows
Normalizing flows (Rezende and Mohamed, 2015; Kobyzev et al., 2020; Papamakarios et al., 2021) are a set of techniques to learn a probability density via a parameterized diffeomorphism $f_{\theta}$ which maps from data to latent space. The probability density is obtained through the change of variables formula:
@@
where $\pi$ is a simple latent distribution. A common strategy to design normalizing flows is through coupling blocks (Dinh et al., 2014, 2016; Durkan et al., 2019), where $f_{\theta}$ is the composition of a series of blocks, each with a triangular Jacobian. This makes the determinant easy to calculate, and hence the above formula is tractable.
@@
4.4 Density-weighted graph algorithm
In order to better estimate probability densities, we use a normalizing flow trained on the data and directly use the density predicted by the model. In order to approximate the geodesic distance between close neighbors, we assume that the geodesic connecting them is approximately a straight line, and estimate its length by a discrete approximation:
$\operatorname{dist}(x_{1},x_{2})\approx\sum_{i=1}^{S}\frac{\|y_{i}-y_{i-1}\|}{p_{\theta}(y_{i-1/2})}$ (11)
Here the $y_{i}$ are points which interpolate $x_{1}$ and $x_{2}$ with $y_{0}=x_{1}$ and $y_{S}=x_{2}$, and $y_{i-1/2}$ is the midpoint of $y_{i-1}$ and $y_{i}$.
This allows us to take a k-nearest neighbors graph of the data and replace the edge weights by the approximate distance. This results in a sparse graph, to which we can apply Dijkstra’s algorithm to estimate shortest paths and distances between any two points. See Algorithm 2.
Algorithm 2 Construction of a density-weighted graph which can subsequently be used to compute shortest paths and distances between points. $X$ is the data matrix, $p$ the probability density, $\beta$ the inverse temperature, $k$ the number of neighbors, and $S$ the number of segments with which to approximate the integral. The computations can also be performed in the log space, returning log edge weights, which is more numerically stable in higher dimensions.
function DensityGraph($X,p,\beta,k,S$)
     $\mathcal{G}\leftarrow\texttt{KNN}(X,k)$triangleright Construct $k$-nearest neighbor graph
     for $(l,m)\in\mathcal{G}$ $\triangleright$ Iterate over edges
         $y_{i}\leftarrow(1-\frac{i}{S})X_{l}+\frac{i}{S}X_{m}$$\triangleright$ For $i=0,\dots,S$
         $y_{i-1/2}\leftarrow\frac{1}{2}(y_{i-1}+y_{i})$$\triangleright$ For $i=1,\dots,S$
         $\mathcal{G}_{lm}\leftarrow\sum_{i=1}^{S}\|y_{i}-y_{i-1}\|/p(y_{i-1/2})^{\beta}$
     end for
     return $\mathcal{G}$
end function
4.5 Relaxation with a learned score model
The score function is the gradient of the log-likelihood with respect to spatial inputs, i.e., $s(x)=\nabla_{x}\log p(x)$. Note the difference from the definition commonly used in statistic, where the derivative is with respect to model parameters.
Score matching (Hyvärinen, 2005) is a family of methods to estimate this function from data, including variants such as sliced score matching (Song et al., 2020a), denoising score matching (Vincent, 2011) and diffusion models (Song et al., 2020b).
@@ 4.5 Relaxation with a learned score model
We can learn the score of $p(x)$ by minimizing the following objective (known as the score matching objective):
@@
Since only the score is needed for the relaxation algorithm described in Section 4.1, we can use a trained score model to perform relaxation to approximate geodesics. In practice, we initialize the path from a graph-based method to speed up convergence and avoid local minima.
@@ 5 Experiments
After solving for a shortest path $\varphi$ between two points $x_{1}$ and $x_{2}$, we compute its length $L(\varphi)$ by using the ground truth density function, then compare this value to the ground truth distance between the two points:
$M(\varphi)=\log\frac{L(\varphi)}{\operatorname{dist}(x_{1},x_{2})}$ (14)
Provided we have access to the ground truth density, this performance metric can be calculated independently of how $\varphi$ is obtained, and has a value of zero only if $\varphi$ is the ground truth geodesic. The metric will otherwise take positive values.
5.1 Scaling the dataset
Figure 3: The length of shortest paths in power-weighted distance graphs do not converge to the ground truth with increasing sample size (green). In contrast, learned and ground-truth density weighted graphs (orange and green) do converge, at similar rates.
In this section we perform experiments on a set of 5 two-dimensional datasets for which we know the ground truth density (see Fig. 3). Details on how the datasets were created can be found in Appendix B.
5.1.1 Power-weighted graph paths do not converge with increasing sample size
We first estimate paths using the power-weighted graph method (Bijral et al., 2012), applied to a k-nearest neighbor graph. We find that in all 5 datasets, increasing the sample size leads to saturating performance (see Fig. 3, green lines). In Fig. 6 in the appendix we show that increasing the number of neighbors does not improve performance.
5.1.2 Poor convergence is due to poor density estimation
In order to investigate the root cause of the poor convergence, we designed other graph methods which use the nearest neighbor estimator for the density, but do not use power-weighted graph edges. We use 4 variants for the density estimate (see Section C.1). We find that all 4 have very similar performance to the power-weighted graph, whereas the equivalent estimators using the ground truth density quickly converge. See Section C.1 for plots supporting this claim.
5.1.3 Learning the density leads to convergent paths
Since the problem lies in poor density estimation, we hope to fix the problem by better density estimation. To this end, we use a normalizing flow to learn the density from samples, then use Algorithm 2 to construct a graph over which we compute shortest paths. We find that the paths obtained converge at close to the same rate as paths obtained by an equivalent method using the ground truth densities (Fig. 3 orange and blue lines). Full experimental details can be found in Appendix B.
@@ 5.2 Scaling the dimension
Figure 4: Mean log error with a sample size of 200,000. Graph-based paths (the first three methods) converge very slowly in higher dimensions with increasingly worse performance as dimension increases. When using the scaled metric ($\beta=1/D$), convergence improves over the unscaled metric, but all graph-based methods perform similarly in higher dimensions. Doing relaxation with a learned score function (red line) leads to much faster convergence, avoiding the curse of dimensionality.
Machine learning datasets often lie in high-dimensional spaces, so it is crucial that any method applied to them scales well, avoiding the curse of dimensionality as much as possible. We test this by taking one of our two-dimensional datasets, namely the standard normal, and scaling it up to 25 dimensions. We perform experiments both for the scaled and unscaled version of the metric.
5.2.1 Graph-based paths do not converge well in higher dimensions
Our first finding is that all graph-based methods show declining performance as the dimension increases, for both unscaled and scaled metrics (Fig. 4). This can be straightforwardly explained by the curse of dimensionality: with increasing dimension but fixed sample size, it is increasingly unlikely that points in the dataset lie close to the true trajectory. The shortest path in the graph becomes increasingly rough in contrast to the smooth ground truth trajectory.
5.2.2 Relaxation using a learned score function improves convergence rates
In order to obtain smoother paths, we turn to the same relaxation algorithm we use to solve for ground truth geodesics. Instead of using the ground truth score function, we use a learned score function. An obvious choice would be to use automatic differentiation to get the score from a trained normalizing flow, but we find the resulting scores are too noisy to be useful, inhibiting reliable convergence of the relaxation scheme. Instead, we learn a score model with sliced score matching, which results in a smoother function suitable for performing relaxation. Full experimental details can be found in Appendix B.
We find that this fixes the convergence problems, both for the unscaled and scaled metrics, though convergence is more reliable in the scaled case (Fig. 4). Please see Fig. 8 and Fig. 9 in the appendix for more detailed evidence.
@@ 6 Conclusion
In conclusion, our study focuses on understanding the convergence properties of various algorithms to ground-truth geodesics by employing toy distributions with known densities. We investigate the standard normal distribution in dimensions higher than 2, highlighting the challenges faced by graph-based methods even in this simple scenario. Our findings demonstrate significant improvements through the use of normalizing flows and score models, yet further research is needed to address more complex distributions and real-world data where the underlying density is unknown. Future work should aim to unify the roles of normalizing flows and score models into a single cohesive model, enhancing the method’s applicability and efficiency. Additionally, a deeper theoretical analysis is required to understand why our proposed methods converge effectively, while previous methods with consistency guarantees do not. This work lays the groundwork for practical applications of density-based distances, setting the stage for more robust and scalable solutions in metric learning.
@@ A.1 Geodesic equation
The geodesic equation which characterizes paths of zero acceleration $\gamma$ is
$\ddot{\gamma}^{i}+\Gamma^{i}_{jk}\dot{\gamma}^{j}\dot{\gamma}^{k}=0$ (15)
where $\Gamma^{i}_{jk}$ are the Christoffel symbols of the second kind
@@
and $g^{il}$ is the inverse of $g$, i.e. such that $g^{il}g_{lj}=\delta^{i}_{j}$.
Our metric has the form $g_{ij}=\lambda^{2}\delta_{ij}$, meaning that
@@
$=\frac{\partial\log\lambda}{\partial x^{k}}\delta^{i}_{j}+\frac{\partial\log\lambda}{\partial x^{j}}\delta^{i}_{k}-\frac{\partial\log\lambda}{\partial x^{l}}\delta^{il}\delta_{jk}$ (19)
In this case $\lambda=p^{-\beta}$ and hence $\frac{\partial\log\lambda}{\partial x}=-\beta\frac{\partial\log p}{\partial x}=-\beta s$, where $s=\frac{\partial\log p}{\partial x}$ is the score function of the probability density. As a result
@@
Substituting into the geodesic equation leads to
@@
which can be written in vector notation as
Let $f:[0,1]\rightarrow[0,1]$ be a strictly increasing diffeomorphism. Given a curve $\gamma:[0,1]\rightarrow\mathbb{R}^{n}$, consider the curve $\varphi:[0,1]\rightarrow\mathbb{R}^{n}$ defined such that $\varphi(u)=\gamma(f^{-1}(u))$ or equivalently $\varphi(f(t))=\gamma(t)$. The derivatives of the two curves are related by
@@
and
[…]
where derivatives of $\varphi$ are with respect to $u=f(t)$. By substituting these expressions into the geodesic equation, we arrive at the equivalent equation for $\varphi$:
$\ddot{\varphi}\dot{f}^{2}+\dot{\varphi}\ddot{f}-2\beta(s(\gamma)\cdot\dot{\varphi}\dot{f})\dot{\varphi}\dot{f}+\beta s(\gamma)\|\dot{\varphi}\dot{f}\|^{2}=0$ (25)
We would like $\varphi$ to have a constant Euclidean speed (the magnitude but not necessarily the direction of the velocity is constant), meaning
@@
Multiplying the geodesic equation by $\dot{\varphi}$, this is equivalent to requiring
$\|\dot{\varphi}\|^{2}\ddot{f}-2\beta(s(\gamma)\cdot\dot{\varphi})\|\dot{\varphi}\|^{2}\dot{f}^{2}+\beta(s(\gamma)\cdot\dot{\varphi})\|\dot{\varphi}\|^{2}\dot{f}^{2}=0$ (27)
implying that
@@
Resubstituting into the geodesic equation and dividing by $\dot{f}^{2}$ leads to the result:

### v2

@@ 2.5 Normalizing Flows
Normalizing flows (Rezende & Mohamed, 2015; Kobyzev et al., 2020; Papamakarios et al., 2021) provide a powerful framework for learning complex probability distributions. The key idea is to transform a simple base distribution (like a standard normal) into the target distribution through a series of invertible transformations. Formally, this is achieved via a parameterized diffeomorphism $f_{\theta}$ which maps from data to latent space. The probability density is obtained through the change of variables formula:
@@
where $\pi$ is a simple latent distribution. A common strategy to design normalizing flows is through coupling blocks (Dinh et al., 2014, 2017; Durkan et al., 2019), where $f_{\theta}$ is the composition of a series of blocks, each with a triangular Jacobian. This makes the determinant easy to calculate, and hence the above formula is tractable.
@@
2.6 Score Matching
While normalizing flows provide a way to estimate the density directly, our geodesic equations actually depend on the score function $s(x)=\nabla_{x}\log p(x)$, which is the gradient of the log-likelihood with respect to spatial inputs. Although one could obtain scores by differentiating the density model, we find that these derived scores are too noisy to be used effectively in practice. We speculate this reflects an inherent trade-off between optimal learning of the log density and its gradient, which we illustrate with simple examples in Section C.3. Therefore, we turn to score matching (Hyvärinen, 2005), which offers a direct way to estimate the score function from data, bypassing the need to estimate the density itself.
This family of methods includes several variants such as sliced score matching (Song et al., 2020), denoising score matching (Vincent, 2011) and diffusion models (Song et al., 2021). For our purposes, we use sliced score matching due to its simplicity and ability to estimate the true score, rather than a noisy score as in denoising score matching or diffusion methods.
@@ 2.6 Score Matching
We can learn the score of $p(x)$ by minimizing the score matching objective:
@@
3 Methods
Figure 5: Illustration of geodesic approximations by graph-based methods (left) and relaxation methods (right) on 200 points sampled from the 2D standard normal. Left: the power-weighted shortest path method (Section 2.4) produces the roughest path. The shortest path from an inverse-density weighted graph using a model density (Section 3.2) is more direct, but not quite as close to the true geodesic as the equivalent method using the ground truth density. The model is defined as Gaussian with the same mean and covariance as the data. Right: relaxation (Algorithm 1) on the model score function results in a very close approximation to the true geodesic.
3.1 Ground Truth Distances through Relaxation
We can solve Equation 4 in several ways to find geodesics between two points. The shooting method starts at one point with an initial velocity guess and integrates the equation forward, iteratively adjusting the initial velocity to reach the target point. Alternatively, the relaxation method discretizes the interval $[0,1]$ into $n$ equal time steps of size $h=1/n$, writing $\varphi_{i}=\varphi(ih)$ for the curve values at these points. Starting from an initial path (either a straight line between endpoints or a more informed guess), it gradually updates the intermediate points while keeping the endpoints fixed to better satisfy the differential equation. Each method has its advantages: shooting is more efficient for exploring geodesics from a starting point, while relaxation is more robust for finding paths between specific endpoints, especially in complex density landscapes where shooting might miss the target. We opt to use relaxation.
Equation 4 leads to the relaxation scheme given in Algorithm 1, which updates each point along the curve based on its neighbors and the local score function. The random ordering of updates helps avoid oscillatory solutions and improves convergence. See Appendix A for a detailed derivation.
3.2 Density-Weighted Graph Algorithm
To improve upon previous graph-based methods, we replace their density estimates with those from a normalizing flow trained on the data. For nearby points, we can approximate the geodesic connecting them as a straight line and estimate its length through numerical integration. Specifically, given two points $x_{1}$ and $x_{2}$, we estimate their distance as:
$\operatorname{dist}(x_{1},x_{2})$ $\approx\sum_{i=1}^{S}\frac{\|y_{i}-y_{i-1}\|}{p_{\theta}(y_{i-1/2})^{\beta}}$ (11)
$=\frac{\|x_{1}-x_{2}\|}{S}\sum_{i=1}^{S}\frac{1}{p_{\theta}(y_{i-1/2})^{\beta}}$ (12)
where $\{y_{i}\}_{i=0}^{S}$ are equally spaced points along the line from $x_{1}$ to $x_{2}$ (i.e., $y_{0}=x_{1}$ and $y_{S}=x_{2}$), and $y_{i-1/2}$ denotes the midpoint between consecutive points.
We use this distance measure to construct a k-nearest neighbors graph, with edge weights given by the approximate geodesic distances rather than Euclidean distances (see Algorithm 2). The resulting sparse graph structure allows us to efficiently compute distances between any pair of points using Dijkstra’s shortest path algorithm. Due to the large dynamic range of the density values involved, we perform all distance calculations and graph operations in log space to maintain numerical stability.
Algorithm 1 One step of relaxation (iterate until convergence)
  Input: trajectory $\varphi=\{\varphi_{i}\}_{i=0}^{n}$, score function $s$, inverse temperature $\beta$
  $I=\{1,\dots,n-1\}$
  $\mathtt{shuffle}(I)$
  for $i$ in $I$ do
     $v_{i}=\frac{1}{2}(\varphi_{i+1}-\varphi_{i-1})$
     $w_{i}=\frac{1}{2}\beta[s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}]$
     $\varphi_{i}=\frac{1}{2}(\varphi_{i+1}+\varphi_{i-1})+w_{i}$
  end for
  Return: $\varphi$
Algorithm 2 Construction of density-weighted graph
  Input: data matrix $X$, density function $p$, inverse temperature $\beta$, neighbors $k$, segments $S$
  $\mathcal{G}=\mathtt{KNN}(X,k)$Construct k-nearest neighbor graph}
  for edge $(l,m)$ in $\mathcal{G}$ do
     $\Delta x=(X_{m}-X_{l})/S$
     for $i=1$ to $S$ do
        $y_{i-1/2}=X_{l}+(i-0.5)\Delta x$
     end for
     $\mathcal{G}_{lm}=\Delta x\sum_{i=1}^{S}1/p(y_{i-1/2})^{\beta}$
  end for
  Return: $\mathcal{G}$
Figure 6: The length of paths in learned and ground-truth density weighted graphs (orange and blue) converge to the ground truth at roughly the same rate with increasing sample size, while the power-weighted graph method (green) converges at a far inferior rate.
3.3 Relaxation with a Learned Score Model
The relaxation algorithm described in Section 3.1 requires only the score function $s(x)$, not the density itself. This observation allows us to use a score model trained directly via score matching, bypassing potential issues with density estimation. In practice, we find this approach more effective than computing scores by differentiating a normalizing flow (see experimental results in Section 4).
To ensure robust convergence, we initialize the path using the graph-based method from the previous section. This initialization provides a reasonable approximation of the geodesic, helping the relaxation process avoid local minima and converge more quickly to the optimal path.
@@ 4 Experiments
We have published the code to reproduce these experiments in the following GitHub repository: https://github.com/vislearn/Fermat-Distance.
@@
To evaluate the quality of computed paths, we introduce the Log Path Ratio (LPR) metric that can be applied to any path $\varphi$ connecting two points $x_{1}$ and $x_{2}$. We compute the path length $L(\varphi)$ using the ground truth density and compare it to the true geodesic distance:
$\operatorname{LPR}(\varphi)=\log\frac{L(\varphi)}{\operatorname{dist}(x_{1},x_{2})}$ (13)
This metric is zero only for the true geodesic and positive otherwise, providing a principled way to evaluate different methods when ground truth densities are available.
4.1 Convergence Analysis on 2D Datasets
To evaluate convergence behavior, we analyze five two-dimensional datasets with known ground truth densities, ranging from simple unimodal distributions to complex multimodal mixtures (shown in Figure 6). For each dataset, we compute the average LPR over 1000 randomly sampled paths. Full experimental details and dataset descriptions are provided in the supplementary material.
4.1.1 Poor Convergence of Power-Weighted Methods
We first examine the power-weighted graph method of Bijral et al. (2011), which constructs paths on k-nearest neighbor graphs. Despite theoretical guarantees, we find that performance improves extremely slowly with increasing sample size across all datasets (Figure 6, green lines).
4.1.2 Density Estimation is the Bottleneck
To identify the source of this poor convergence, we developed several variant methods that use different nearest-neighbor density estimators while maintaining the same graph structure (see supplementary material for details). All variants show similarly poor performance, while their counterparts using ground truth densities converge rapidly. This clearly identifies density estimation as the primary bottleneck.
4.1.3 Improved Convergence through Better Density Estimation
Having identified poor density estimation as the core issue, we replace nearest-neighbor estimates with a learned normalizing flow model. Using this density estimate in Algorithm 2, we find that the resulting paths converge at nearly the same rate as those computed using ground truth densities (Figure 6, orange and blue lines). This demonstrates that modern density estimation techniques can effectively close the gap between theoretical guarantees and practical performance. Full experimental details can be found in the supplementary material.
@@ 4.2 Scaling the Dimension
Figure 7: Mean LPR with a sample size of 200,000. Graph-based methods (first three) show rapidly degrading performance as dimension increases, even with the scaled metric ($\beta=1/D$). In contrast, relaxation with a learned score function (red line) maintains good performance across dimensions, effectively avoiding the curse of dimensionality.
Machine learning datasets often lie in high-dimensional spaces, making dimensional scaling behavior crucial for practical applications. We investigate this by extending one of our two-dimensional datasets (the standard normal) up to 25 dimensions, testing both the scaled ($\beta=1/D$) and unscaled ($\beta=1$) versions of the metric.
4.2.1 Graph-Based Methods Fail in Higher Dimensions
All graph-based methods show rapidly degrading performance as dimension increases, regardless of whether we use scaled or unscaled metrics (Figure 7). This failure is a direct consequence of the curse of dimensionality: with fixed sample size but increasing dimension, the probability of finding sample points near the true geodesic becomes vanishingly small. As a result, graph-based paths become increasingly jagged approximations of the smooth ground truth trajectory.
4.2.2 Score-Based Relaxation Overcomes Dimensional Scaling
To address this fundamental limitation of graph-based methods, we apply our relaxation algorithm (Algorithm 1) with a learned score function. While one might expect to obtain scores by differentiating a trained normalizing flow, we find these derived scores too noisy for reliable convergence (see Section C.3 for analysis of this phenomenon). Instead, direct score estimation via sliced score matching provides smoother gradients that enable stable relaxation.
This score-based approach dramatically improves performance, maintaining good convergence across dimensions for both scaled and unscaled metrics, though the scaled version ($\beta=1/D$) shows more reliable behavior (Figure 7). Detailed experimental results and analysis can be found in the supplementary material.
4.3 MNIST
We also perform preliminary experiments on MNIST within the latent space of an autoencoder, as described in Section C.1. We find that we obtain interpretable and reasonable geodesics connecting digits within this space, and that digits cluster together in an unsupervised way.
@@ 5 Conclusion
We have presented a systematic study of geodesic computation in density-based metrics, using distributions with known ground truth to rigorously evaluate different approaches. Our investigation reveals several key insights:
First, we demonstrate that previous graph-based methods, despite their theoretical guarantees, converge impractically slowly even for simple distributions like the standard normal. Second, we show that this poor convergence stems primarily from inadequate density estimation rather than the graph approximation itself. Third, we establish that modern deep learning approaches—normalizing flows for density estimation and score matching for gradients—can dramatically improve performance.
However, significant challenges remain. While our methods work well on toy distributions with known densities, extending them to complex real-world data requires further research. Key directions for future work include:
  • Unifying normalizing flows and score models into a single framework to improve efficiency
  • Developing theoretical understanding of why our methods converge effectively while previous approaches with consistency guarantees do not
  • Extending these methods to higher-dimensional densities typical in real applications
This work establishes a foundation for practical density-based distances, bringing theoretical elegance closer to practical applicability in metric learning.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
This work is supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2181/1 - 390900948 (the Heidelberg STRUCTURES Cluster of Excellence). The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.
@@ A.1 Geodesic Equation
We begin by deriving the geodesic equation for our conformal metric. This is a well-known result in differential geometry (see e.g., Carroll (2004, Appendix G)). The key steps are computing the Christoffel symbols and then simplifying the resulting equation.
Lemma A.1 (Christoffel Symbols for Conformal Metrics).
For a conformal metric of the form $g_{ij}=\lambda^{2}\delta_{ij}$, the Christoffel symbols are:
$\Gamma^{i}_{jk}=\frac{\partial\log\lambda}{\partial x^{k}}\delta^{i}_{j}+\frac{\partial\log\lambda}{\partial x^{j}}\delta^{i}_{k}-\frac{\partial\log\lambda}{\partial x^{l}}\delta^{il}\delta_{jk}$ (14)
Proof.
The Christoffel symbols are given by:
@@
For our metric:
@@
The result follows by simplifying the Kronecker delta products. ∎
Theorem A.2 (Geodesic Equation).
For a metric of the form $g_{ij}=p^{-2\beta}\delta_{ij}$, where $p$ is a probability density and $\beta$ is a constant, the geodesic equation in vector notation is:
where $s=\nabla\log p$ is the score function of the probability density.
Proof.
For our metric, $\lambda=p^{-\beta}$ and hence $\frac{\partial\log\lambda}{\partial x}=-\beta\frac{\partial\log p}{\partial x}=-\beta s$. Substituting into the formula for Christoffel symbols:
@@
The geodesic equation $\ddot{\gamma}^{i}+\Gamma^{i}_{jk}\dot{\gamma}^{j}\dot{\gamma}^{k}=0$ then becomes:
@@
which simplifies to the desired vector form. ∎
Theorem A.3 (Constant Euclidean Speed Parameterization).
Under a reparameterization that maintains constant Euclidean speed, the geodesic equation becomes:
Proof.
Let $f:[0,1]\rightarrow[0,1]$ be a strictly increasing diffeomorphism and define $\varphi(u)=\gamma(f^{-1}(u))$. The derivatives are related by:
@@
$\ddot{\varphi}\dot{f}^{2}+\dot{\varphi}\ddot{f}$ $=\ddot{\gamma}$ (23)
For constant Euclidean speed, we require:
@@
Substituting into the geodesic equation and solving for $\ddot{f}$ yields:
@@
The result follows by substituting back and simplifying. ∎